OpenStudy (anonymous):

Solve: −2(x − 8) = −10

OpenStudy (katrinakaif):

Distribute.

OpenStudy (katrinakaif):

-2x + 16 = -10

OpenStudy (katrinakaif):

Bring the constants to one side.

OpenStudy (katrinakaif):

-2x = -10 - 16 -2x = -26

OpenStudy (katrinakaif):

Divide by -2 on both sides.

OpenStudy (katrinakaif):

x = 13

OpenStudy (katrinakaif):

A medal would be appreciated ;)